Grade 12th passAlgebra

Find the equations of the straight lines which pass through the intersection of 3X-4Y+1=0and 5X+Y=1 and which cut off equal intercepts from the axes

guru
9 Years agoGrade 12th pass

1 Answer

Amreen
9 Years ago
Solve the two eqns
3x-4y=-1
(5x+y=1)4
after solving, x=3/23
y=8/23
 
a=b(given)
 
so,usng the eqn
x/a+y/b=1
solve for a by substituting values of x and y
3/23a+8/23a=1
:a=11/23
 
eqn is 23x+23y=11